I posted this problem before but my other E-mail address is actting up please send any info on my Question to this address. Sorry and Thank you.


Could you show me the steps on how to solve these problems I don’t need to see the final answer just how you got it.


3 trains—1 eastbound, 1westbound, and 1northbound—leave a city at the same time. The speed of the northbound train is 20 mph greater than the speed of the eastbound train After 7 hours, the distance between the westbound train and the eastbound train is 700 miles. Twice the speed of the westbound train is 90 mph more than the speed of the northbound train. . Find the speed of the 3 trains.
I know I didn’t set up the equations right but I did the best I could
3 trains—1 eastbound, 1westbound, and 1northbound—leave a city at the same time.
Eastbound = X, westbound = Y, northbound = Z.

The speed of the northbound train is 20 mph greater than the speed of the eastbound train
20 >X=Z

After 7 hours, the distance between the westbound train and the eastbound train is 700 miles.
Y-X=700 miles

Twice the speed of the westbound train is 90 mph more than the speed of the northbound train.
2Y=90+Z
Find the speed of the 3 trains. I have no idea how to do that.
Thank you Very Much

Answer by Paul(988)

Eastbound train's speed = x
Northbound train's speed = x+20
Westbound train's speed = y
Two varible equations so two statements are given.
In first statement:
7(y+x)=700
7y+7x=700
y=100-x
So speed of the westbound train is >100 of the eastbound train.
Second statment:
2y=90+(x+20)
2y=110+x
Subsitutue for y:
2(100-x)=110+x
200-2x=110+x
-3x=-90
x=30
30+20=50 (northbound)
y=100-30
y=70 (westbound)
Hence, the speed of the eastbound train is 30mph, the speed of northbound train is 50mph, and the speed of the westbound train is 70mph.
Paul.